OpenStudy (anonymous):

Given \(a_1 = 1\), \(b_1=-\frac{1}{2}\), and \(a_n = \frac{a_{n-1}}{b_{n-1}}\) and \(b_n = a_{n-1}*b_{n-1}\) for all n>1. Show that \((a_n)_{n\in\mathbb{N}}\) and \((b_n)_{n\in\mathbb{N}}\) diverge.

OpenStudy (experimentx):

\[ a_n\cdot b_n = \frac{a_{n-1}}{b_{n-1}}\cdot a_{n-1}b_{n-1} = (a_{n-1})^2 \] similarly, \[ \frac{a_n}{ b_n} = \frac{a_{n-1}}{b_{n-1}}\cdot \frac 1{ a_{n-1}b_{n-1}} = \frac 1 {b_{n-1}^2}\]

OpenStudy (experimentx):

\[ a_n = \frac{a_{n-1}}{b_{n-1}} = \frac{1}{b_{n-1}^2}\] and \[ b_n =a_{n-1}b_{n-1} = a_{n-2}^2 \]

OpenStudy (experimentx):

* \[ a_n = \frac{a_{n-1}}{b_{n-1}} = \frac{1}{b_{n-2}^2} \]

OpenStudy (experimentx):

\[ a_{n} = \frac{1}{(a_{n-4})^4}\]

OpenStudy (experimentx):

\[ a_2 = -2, a_6 = 1/16, a_{10} = 16^4 \dots \]

OpenStudy (experimentx):

\[ a_{18} = (-2)^{4^3} \\ \huge {a_{8n+2} = (-2)^{4^{(2n)}}}\] easy to see that it diverges as n->infty

OpenStudy (experimentx):

similarly it can be shown via recursion relation for b_n GTG see you later

OpenStudy (anonymous):

Hmm.... So, I can show the subsequence of it diverges and claim that it diverges by the comparison test!!!

OpenStudy (experimentx):

yep

OpenStudy (experimentx):

a sequence converges if and only if all it's sub-sequence converges ... or you can use the cauchy sequence test too.

OpenStudy (experimentx):

if particular if \( a_{f(n)} \) is sub-sequence of \( a_n \) then \( a_n \) converges if \( a_{f(n)} \) converges for all \(f:\Bbb N \to \Bbb N \).

OpenStudy (experimentx):

*iff